# Metastability induced by non-reciprocal adaptive couplings in Kuramoto models

### v1

@@ III.5 Impact of disorder via noise or natural frequencies
In Kuramoto models with two non-reciprocally interacting populations, disorder in the frequencies or noise can stabilize a chiral phase, which is chiral in the sense that instead of static phase alignment a chiral motion is observed such that in spite of zero natural frequencies, oscillations spontaneously rotate either clockwise or anti-clockwise [23]. In our version of the Kuramoto model with one population and adaptive non-reciprocal couplings to all other oscillators, the role of disorder or noise that we observed is toward an increase of escapes from a two anti-phase synchronized cluster state. For the strongest noise amplitude in Fig.11 (d), the order parameter $R_{2}$ shows larger deviations from $1$, indicating that temporarily several clusters have formed, but $R_{2}$ remains close to $1$ and returns to this value again and again, so that still remnants of the anti-phase synchronized two clusters are still visible. Since the phase difference between the two clusters is roughly $\pi$, noise is in general not sufficient for an escape where many oscillators overcome the distance and change their cluster affiliation at the same time, just a few succeed as the panels (c) in Fig. 9 and 10 show. In particular, noise does not improve synchronization at an intermediate noise level; due to a large phase difference, it monotonically weakens the perfect anti-phase cluster synchronization, allowing for longer escapes, and broadens a bit the in-phase synchronization within a cluster.
@@ III.6 Effects of a sparse network topology
In the next version of the model inspired by the Ref. [12], the dynamics of the oscillators and their adaptive couplings have been modified to account for a connected but sparse random network topology. The coupling matrix $k_{ij}$ now evolves only along the links of a network represented by the adjacency matrix $A=[A_{ij}]$. The phase dynamics are therefore weighted by the network structure:
@@ III.7 Kuramoto Model with non-reciprocal and adaptive frequencies
The time evolution of the order parameters and upper and lower couplings looks qualitatively the same as before. Whenever $R_{2}$ approaches $1$, the oscillators’ phases naturally divide into two groups. Despite all oscillators starting with $\omega_{i}=0$, they begin to oscillate due to the nonzero value of $\Gamma_{0}=0.1$.
We show a snapshot at $t=15000$ of the phases and frequencies $\tilde{\omega_{i}}$ in Fig. (13) (a) and (b), where each oscillator has a distinct $\tilde{\omega_{i}}$ that oscillates over time. The 10 oscillators are divided into two groups separated by a phase difference of $\pi$, although the frequencies are quite heterogeneous. In panel (c), we display a snapshot of the coupling matrix $k_{ij}$. The lower triangle of the matrix does not yet converge to its equilibrium values of $\pm 1$ due to the slow adaptation rate $\varepsilon_{2}=0.0001$, the upper one does. One can easily read off which oscillators remain in the same cluster with upper couplings being attractive and which ones are in opposite clusters with upper couplings being repulsive. The long-time evolution of the phases $\theta_{i}$ (Fig. (13) (e)) indicates how the anti-phase synchronization patterns of phases change together with the frequency patterns (d) in time intervals between abrupt changes of the slope of $\tilde{\omega}_{i}$ and $\theta_{i}$. Note that $\tilde{\omega}_{i}$ play the role of effective natural frequencies. Obviously they can considerably differ from the actual frequencies $\dot{\theta}_{i}$, being larger in the time interval between $t\in[2000,4000]$ than in the one between $t\in[8000,10000]$, although the opposite order is seen for $\tilde{\omega}_{i}$.
@@ IV Conclusions and Outlook
We have studied generalized Kuramoto models of classical oscillators which have applications in particular to model oscillatory dynamics in neuronal networks. Strengths or weights assigned to synapses are in general not symmetric, therefore it is natural to model their dynamics via non-reciprocal couplings, where we have chosen Hebbian and anti-Hebbian rules for adaptation. Moreover, the time scales on which adaptation takes place need not be the same in both directions. For a certain range of parameters and in this combination of asymmetric coupling‘s type and time-scale, we observe metastability of (in general) two anti-phase clusters of synchronized oscillators where a few oscillators (“switcher oscillators”) change their cluster affiliation. This is induced by a dynamical switch of the coupling type (attractive or repulsive) of connections $ij$ to nodes $j$ from swing oscillators $i$, which happen on different time scales, depending on whether $i<j$ or $i>j$. We have seen temporary remnants of dynamical realizations of Harari’s theorem, that is, the existence of two anti-phase synchronized clusters on an all-to-all topology, as long as the dynamics on the second time scale, parameterized by $1/\epsilon_{2}$, can be neglected.
In view of brain dynamics it is of interest to identify interactions which are Hebbian-like in one direction and anti-Hebbian-like in the opposite direction together with oscillatory subunits which change the clusters of synchronized units they belong to. For future work it would be very interesting to search for dynamical phases as result of non-reciprocal interactions in models of limit-cycle oscillators such as Stuart-Landau or Fitz-Hugh Nagumo oscillators which are also relevant as models for brain dynamics. Otherwise there are many ways to implement non-reciprocity to formerly symmetric couplings, for example in multiplex networks [37, 38], in networks with designed two anti-phase clusters [39, 40], or networks leading to emergent anti-phase clusters such as in [41, 42], or in networks with geometric frustration which have been constructed such that a continuum of attractors was generated together with a (dynamically generated) Watanabe-Strogatz phenomenon of dimensional reduction [43, 44, 36].
@@ Appendix A Appendix: Stability analysis of two anti-phase clusters of arbitrary sizes
Consider the original fully connected Kuramoto model in the absence of noise ($\sigma_{\mathrm{noise}}=0$) with $\omega_{i}=0$ $\forall$ $i=1,2,3,\cdots,N$. Assume the oscillators form two anti-phase clusters, named Cluster A and Cluster B, of sizes $N_{A}$ and $N_{B}=N-N_{A}$, with phases
@@ Acknowledgments
The original contributions presented in the study are included in the article/Supplementary Material. Further inquiries can be directed to the corresponding author.

### v2

@@ III.5 Impact of disorder via noise or natural frequencies
In Kuramoto models with two non-reciprocally interacting populations, disorder in the frequencies or noise can stabilize a chiral phase, which is chiral in the sense that instead of static phase alignment a chiral motion is observed such that in spite of zero natural frequencies, oscillators spontaneously rotate either clockwise or anti-clockwise [33]. In our version of the Kuramoto model with one population and adaptive non-reciprocal couplings to all other oscillators, the role of disorder or noise that we observed is toward an increase of escapes from a two anti-phase synchronized cluster state. For the strongest noise amplitude in Fig.9 (d), the order parameter $R_{2}$ shows larger deviations from $1$, indicating that temporarily several clusters have formed, but $R_{2}$ remains close to $1$ and returns to this value again and again, so that remnants of the anti-phase synchronized two clusters are still visible. Since the phase difference between the two clusters is roughly $\pi$, noise is in general not sufficient for an escape where many oscillators overcome the distance and change their cluster affiliation at the same time, just a few succeed as the panels (c) in Fig. (8) show. In particular, noise does not improve synchronization at an intermediate noise level; instead, due to a large phase difference, it monotonically weakens the perfect anti-phase cluster synchronization, allowing for longer escapes, and broadens a bit the in-phase synchronization within a cluster.
@@ III.6 Effects of a sparse network topology
In the next version of the model, inspired by the Ref. [21], the dynamics of the oscillators and their adaptive couplings are modified to account for a connected but sparse random network topology. The coupling matrix $k_{ij}$ now evolves only along the links of a network represented by the adjacency matrix $A=[A_{ij}]$. The phase dynamics are therefore weighted by the network structure:
@@ III.7 Kuramoto Model with non-reciprocal and adaptive frequencies
The time evolution of the order parameters and upper and lower couplings looks qualitatively the same as before. Whenever $R_{2}$ approaches $1$ (not displayed in Fig. (11)), the oscillators’ phases naturally divide into two groups. Despite all oscillators starting with $\omega_{i}=0$, they begin to oscillate due to the nonzero value of $\Gamma_{0}=0.1$.
We show a snapshot at $t=15000$ of the phases and frequencies $\tilde{\omega_{i}}$ in Fig. (11) (a) and (b), where each oscillator has a distinct $\tilde{\omega_{i}}$ that oscillates over time. The 10 oscillators are divided into two groups separated by a phase difference of $\pi$, although the frequencies are quite heterogeneous. In panel (c), we display a snapshot of the coupling matrix $k_{ij}$. The lower triangle of the matrix does not yet converge to its equilibrium values of $\pm 1$ due to the slow adaptation rate $\varepsilon_{2}=0.0001$, the upper one does. One can easily read off which oscillators remain in the same cluster with upper couplings being attractive and which ones are in opposite clusters with upper couplings being repulsive. The long-time evolution of the phases $\theta_{i}$ (Fig. (11) (e)) indicates how the anti-phase synchronized clusters change together with the frequency patterns (d) in time intervals between abrupt changes of the slope of $\tilde{\omega}_{i}$ and $\theta_{i}$. Note that $\tilde{\omega}_{i}$ play the role of effective natural frequencies. Obviously they can considerably differ from the actual frequencies $\dot{\theta}_{i}$, being larger in the time interval between $t\in[2000,4000]$ than in the one between $t\in[8000,10000]$, although the opposite order is seen for $\tilde{\omega}_{i}$.
@@ IV Conclusions and Outlook
We have studied generalized Kuramoto models of classical oscillators which have applications in particular to model oscillatory dynamics in neuronal networks. Strengths or weights assigned to synapses are in general not symmetric, therefore it is natural to model their dynamics via non-reciprocal couplings, where we have chosen Hebbian and anti-Hebbian rules for plasticity. Moreover, the time scales on which adaptation takes place need not be the same in both directions. For a certain range of parameters and in this combination of asymmetric coupling’s type and time-scale, we observe metastability of (in general) two anti-phase clusters of synchronized oscillators where a few oscillators (“switcher oscillators”) change their cluster affiliation. This is induced by a dynamical switch of the coupling type (attractive or repulsive) of connections $ij$ to nodes $j$ from switcher oscillators $i$, which happen on different time scales, depending on whether $i<j$ or $i>j$. We see temporary remnants of dynamical realizations of Harary’s theorem, that is, the existence of two anti-phase synchronized clusters on an all-to-all topology, as long as the dynamics on the second time scale, parameterized by $1/\epsilon_{2}$, can be neglected.
Regarding the system size, individual oscillators may represent mesoscopic neural populations such as cortical columns [51]. Therefore a small system size of $N=10$ or $N=20$ with pronounced metastability would include many neurons; here the increasing instability for large $N$, derived from our scaling analysis, would not apply to the individual neurons, but to a large number of brain areas. The ‘switcher’ mechanism would mean that some areas are flexible with respect to the choice of the cluster with which to synchronize. In view of experiments, first steps may amount to identify interactions which are Hebbian-like in one direction and anti-Hebbian-like in the opposite direction together with the time scales they are acting on.
In our current model with non-reciprocal phase shifts of $\pm\pi/2$, the system favors two anti-phase clusters. Different options exist to obtain states with more than two clusters or different collective behaviors (e.g., a splay state [52, 53], or generalized chimeras [54, 55]). One would typically need to modify the interaction function to include higher-order interactions [56] or alter the phase lag parameters [20] away from strict orthogonality, as shown in previous studies on clustering in coupled oscillator networks [57, 58]. An open question is as to whether switchers between more than two clusters could be created due to the same kind of non-reciprocal adaptive couplings. Finally, in order to see new collective phenomena in neural networks, one may search for dynamical phases as result of non-reciprocal interactions in models of limit-cycle oscillators such as Stuart-Landau or Fitz-Hugh Nagumo oscillators, including amplitudes apart from oscillator phases as considered in this work.
@@ Appendix A Appendix: Stability analysis of two anti-phase clusters of arbitrary sizes
Consider the original fully connected Kuramoto model (1) in the absence of noise ($\sigma_{\mathrm{noise}}=0$) with $\omega_{i}=0$ $\forall$ $i=1,2,3,\cdots,N$. Assume the oscillators form two anti-phase clusters, named Cluster A and Cluster B, of sizes $N_{A}$ and $N_{B}=N-N_{A}$, with phases
@@ Acknowledgments
The original contributions presented in the study are included in the article. The code used to generate the key findings and animations are publicly available in the GitHub repository [48].